Reports 1-1 of 1 Clear search Modify search
MIF (ITF Control)
takafumi.ushiba - 18:49 Sunday 31 July 2022 (21659) Print this report
ALS DARM was locked

[Yamamoto, Ushiba]


ALS DARM was locked and lasted for more than 2 hours!!

We designed feedback filters from ALS DARM signals to ETMX.
During O3, we used only TM and MN for mass feedback but in this time, we used all three stages:TM, IM, and MN.
Since we haven't measured OLTF, we don't know UGF, cross over frequencies, and so on but designed values are as follows:

UGF: 20 Hz
COF: MN-IM: 3 Hz

Figure 1 shows the spectra of ALS DARM signals with (red) and without (blue) mass feedback.
RMS becomes more than 10 times smaller.

We have already implemented ALS DARM lock in LSC_LOCK guardian and can lock it by requesting ALS_DARM_LOCKED state for the guardian, if all mirrors' alignments are good.

Red spectrum is smaler than blue one between 10 Hz and 100 Hz but this is not due to DARM feedback.
We are not sure the source but somehow spectrum at these frequency regions are breathing.

Images attached to this report
Comments to this report:
takafumi.ushiba - 19:19 Sunday 31 July 2022 (21660) Print this report

We measured OLTF of current ALS DARM loop (fig1).
UGF is about 10 Hz with a phase margin of 40 Hz, which is very close to what we expected.

Images attached to this comment
tomotada.akutsu - 21:27 Sunday 31 July 2022 (21661) Print this report

Bravo oh bravo!

masayuki.nakano - 14:56 Monday 08 August 2022 (21729) Print this report

OMG!! I missed a really great news!!!

I was really surprised by your speed to succeed in the type-A control, and I'm pretty sure that you will lock FPMI and be ready for the noise hunting!!

Go KAGRA, go!!

Search Help